This is not from a textbook.
Write a quadratic equation in the variable x having the given numbers as solutions. Type the equation in standard form, ax^2 + bx + c = 0
solution: 7, only solution
Equation is 0 = ___________
I'm just starting to be able to understand how to compute these kind of equations so to do it backwards is confusing me totally.

Answer by stanbon(75887)

If 7 is the only solution, it has multiplicity two.
Therefore it has two factors that look like (x-7)
EQUATION:
y = (x-7)^2 = x^2-14x+49
=============
Cheers,
Stan H.